OpenStudy (anonymous):

A student has two loans totaled $45,000. The interest of one loan is 5.1% and the interest of the other loan is 3.8%. After one year, the student has paid $1953.10 in interest. What was the amount of each loan?

OpenStudy (anonymous):

0.051 x + 0.038y = 1953.1 x + y = 45000 solve for x and y

OpenStudy (anonymous):

I am not sure how

OpenStudy (anonymous):

ok multiply the 2nd equn, by 0.051 to eliminate x 0.051x + 0.051y = 2295 now subtract this from first equn. -0.013y = - 341.9 y = -341.9 / -0.013 = $26300 - is the amount of first loan x = 45000 - 26300

OpenStudy (anonymous):

ill let you do the last subtraction

OpenStudy (anonymous):

ok this may take a minute

OpenStudy (anonymous):

I am kind of confused....what did you multiply 0.051 by?

OpenStudy (anonymous):

the equation x + y = 45000 i multiplied each of the three terms by 0.051

OpenStudy (anonymous):

OOh okay, after subtracting I got 18,700

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

sweet so the 3.8 loan is 18700 and the 5.1 loan is 26300??

OpenStudy (anonymous):

right

OpenStudy (anonymous):

sweet thank you >D

OpenStudy (anonymous):

ok i gotta go